Is the answer to this equation no solution: 0.5(x - 3) + (1.5 - x) = 5x
Answer by Alan3354(69443)

Is the answer to this equation no solution: 0.5(x - 3) + (1.5 - x) = 5x
----------
0.5(x - 3) + (1.5 - x) = 5x
(x - 3) + 2(1.5 - x) = 10x
x - 3 + 3 - 2x = 10x
x - 3 + 3 - 2x = 10x
-x = 10x
0 = 11x
x = 0